OpenStudy (anonymous):

Your piggy bank has a total of 46 coins in it; some are dimes and some are quarters. If you have a total of $7.00, how many quarters and how many dimes do you have?

OpenStudy (anonymous):

@whpalmer4

OpenStudy (nurali):

25q + 10(46 - q) = 700 5q + 2(46 - q) = 140 5q + 92 - 2q = 140 3q = 48 q = 16 No. of dimes: 46 - 16 30 16 quarters & 30 dimes

OpenStudy (whpalmer4):

@Nurali's work is correct; I personally would have let you do some of the work and some more of the learning, too! :-)

OpenStudy (anonymous):

Thank You!!! @whpalmer4 @Nurali

OpenStudy (nurali):

My Pleasure.